OpenStudy (anonymous):

Find a polynomial for the perimeter and for the area. I have a square box with the top being n+3 and the right side n what is the perimeter ? what is the area?

OpenStudy (trancenova):

perimeter=2n+3 area=n(n+c3)=n^2+3n

OpenStudy (anonymous):

where did c comes from?

OpenStudy (trancenova):

oh sorry, typo!

OpenStudy (anonymous):

She says she has a square box ... O o

OpenStudy (anonymous):

How could it be a square box if sides are not equal.....and box is 3d so we have surface area .... and volume

OpenStudy (trancenova):

hmmm I was thinking that... I just assumed it meant rectangle